OpenStudy (anonymous):

solve each equation: 1/2+ x/6= 18/x

OpenStudy (callisto):

\[\frac{1}{2}+ \frac{x}{6}= \frac{18}{x}\]\[6x(\frac{1}{2}+ \frac{x}{6})= 6x(\frac{18}{x})\]\[3x+x^2 = 108\]\[x^2+3x-108 = 0\]\[(x+12)(x-9) = 0\]Can you solve it ?

OpenStudy (anonymous):

yes